# Long-range Ising model for regional-scale seismic risk analysis

### v1

@@ Abstract
This study introduces the long-range Ising model from statistical mechanics to the Performance-Based Earthquake Engineering (PBEE) framework for regional seismic damage analysis. The application of the PBEE framework at a regional scale entails estimating the damage states of numerous structures, typically performed using fragility function-based stochastic simulations. However, these simulations often assume independence or employ simplistic dependency models among capacities of structures, leading to significant misrepresentation of risk. The Ising model addresses this issue by converting the available information on binary damage states (safe or failure) into a joint probability mass function, leveraging the principle of maximum entropy. The Ising model offers two main benefits: (1) it requires only the first- and second-order cross-moments, enabling seamless integration with the existing PBEE framework, and (2) it provides meaningful physical interpretations of the model parameters, facilitating the uncovering of insights not apparent from data. To demonstrate the proposed method, we applied the Ising model to $156$ buildings in Antakya, Turkey, using post-hazard damage evaluation data, and to $182$ buildings in Pacific Heights, San Francisco, using simulated data from the Regional Resilience Determination (R2D) tool. In both instances, the Ising model accurately reproduces the provided information and generates meaningful insights into regional damages. The study also investigates the change in Ising model parameters under varying earthquake magnitudes, along with the mean-field approximation, further facilitating the applicability of the proposed approach.
@@ 1 Introduction
The Performance-Based Earthquake Engineering (PBEE) framework has been extensively developed and utilized for over 20 years, providing a systematic method to calculate quantifiable metrics for seismic design and post-hazard decision-making [1, 2]. To incorporate the various uncertainties that propagate from ground motions to the seismic performance of a structure, the PBEE framework considers four groups of random variables: (1) ground motion intensity measure (IM), (2) engineering demand parameter (EDP), (3) damage measure or damage state (DS), and (4) loss or decision variable (DV), each described by its probability distribution. Given an earthquake scenario, IM is typically described as a probability distribution $P(IM)$, or as a mean annual probability of exceedance $\lambda(IM)$, assuming a Poisson process model for the temporal occurrence of earthquakes. This intensity measure model is followed by the conditional distribution of EDP given IM, denoted by $P(EDP|IM)$. Subsequently, the conditional distribution of DS given EDP, $P(DS|EDP)$, is established, which can also be expressed as a complementary cumulative distribution function, or component-level fragility function, $G(DS|EDP)$. Finally, $G(DS|EDP)$ is followed by the distribution of DV given DS, $P(DV|DS)$. To obtain the exceedance probability $\lambda(DV)$ of a decision variable, the probability density functions $P(\cdot)$ are often transformed to complementary cumulative distribution functions $G(\cdot)$, yielding the integral:
[…]
@@
Recently, the application of the PBEE framework has been actively extended to a regional scale [7, 8, 9, 10]. In this context, Eq. (1) is typically solved using Monte Carlo methods that simulate regional responses to hazards, hereinafter referred to as regional seismic simulations. This simulation approach is favored because a system-level decision variable may depend on the damage states of all components, rendering Eq. (1) high-dimensional. Similar to the component-level applications, a regional-scale PBEE analysis can be decomposed into four modules: (1) IM generation, (2) EDP estimation, (3) DS evaluation, and (4) DV assessment. For each of these modules, physics-based or statistical methods/models can be employed, offering a spectrum of variations. For instance, to generate IM samples, one can simulate seismic wave propagation [11] to create synthetic seismograms or employ empirical ground motion prediction equations [12]. Nonlinear structural dynamic analysis can be performed to generate EDP samples [10], while several regression models directly connecting EDP to IM are also available [13, 14]. The DS can be obtained from structural analysis or inferred from IM or EDP samples using DS-IM or DS-EDP fragility curves. Lastly, DV can be assessed using loss curves that relate DS or EDP to DV [15], or physics-based simulations [9].
Physics-based simulations can accurately capture the complex interactions and correlations among structural responses and functionalities. However, they are often restricted to scenario-based analyses due to their extensive computational costs. On the other hand, statistical approaches, such as fragility function-based simulations of DS given EDP or IM [16], are computationally feasible but require many ad-hoc, sometimes subjective, assumptions about statistical models. Specifically, applying structural-level fragility functions to regional-scale analysis entails constructing a fragility field that summarizes the joint information across the region. To capture the spatial dependency among structural damages, the fragility field demands additional assumptions and parameters than individual fragility curves. The common practice of using individual fragility functions under the assumption of conditional independence (conditional on site-specific intensity measures) can significantly underestimate system-level failure probabilities [17].
@@
Section 2 overviews the current approaches for evaluating structural damages in the regional-scale PBEE. Subsequently, Section 3 introduces the formulation and parameter estimation of the Ising model in the context of regional seismic analysis. This section also develops the mean-field approximation to qualitatively explain the Ising model parameters. Section 4 offers interpretations of the Ising model parameters in the language of regional seismic risk analysis and explores their relationship with the mean and correlation coefficients of the damage states of structures. Section 5 demonstrates the validity and engineering relevance of the Ising model through two examples using real-world post-hazard survey data and synthetic simulation data, emphasizing the insights derived from the Ising model. Finally, concluding remarks and future research directions are presented in Section 6.
@@ 2 Current approaches for probabilistic structural damage predictions on a regional-scale
For an individual structure, the following fragility function is commonly used to predict its damage state:
@@
The missing information in the above procedure is the correlation between damage states contributed by the similarities of structures. Bazzurro and Luco Nicolas [22] emphasized that accounting for the similarity of structures represents a key distinction between analyzing a single structure and a portfolio of structures in the seismic loss estimation. Similarities between structures, such as the number of stories, construction materials, and structural type, are particularly pronounced in residential neighborhoods where structures share similar designs and are built following the same design codes. This contributes to the correlation between their damage states [17], in addition to the contribution from the spatial coherency of ground motions. To model the structure-to-structure damage correlation, Lee and Kiremidjian [23] presented a simplified approach by assuming equi-correlation among all structures. The study also concluded that the challenge of modeling the damage correlation lies in approximating the joint discrete damage distribution. Due to the difficulty of defining a joint distribution for discrete damage states, Baker and Cornell [24] proposed using the joint continuous distribution for the decision variables given the engineering demand parameters, $f(\boldsymbol{DV}|\boldsymbol{EDP})$, by integrating $f(\boldsymbol{DM}|\boldsymbol{EDP})$ and $f(\boldsymbol{DV}|\boldsymbol{DM})$, thereby bypassing the need for a joint distribution of discrete damage states. An alternative approach, as detailed in [25, 26], models fragility using continuous demand and capacity. This method has become popular because of its convenience.
@@ 3 Long-range Ising model as a joint probability mass function for damage states
The Ising model was originally conceived as a mathematical model for ferromagnetism in statistical mechanics [28, 29]. A key feature of the model is that it represents the maximum entropy distribution constrained by the means and pairwise correlations, making it applicable to a wide variety of fields beyond statistical mechanics [30, 31, 18]. Provided with the first- and second-order cross-moments of the component states in a system, the Ising model is the least structured joint probability distribution for the component states that reproduces the prior information. In this section, we first formulate the Ising model within the context of regional seismic analysis, followed by a discussion on its parameter estimation. Finally, we present the mean-field approximation, which offers a qualitative understanding of the relationship between the Ising model parameters and the given statistics.
@@ 3.1 Formulation
Consider a region consisting of $n$ structures under an earthquake and assume that each structure can be in one of the binary states of safe or failure. Here, “safe” and “failure” are relative terms with respect to a target performance level. The group of structures has $N=2^{n}$ configurations, where the probability of the $I$-th configuration is expressed as:
@@
The entropy $S$ for $\boldsymbol{X}=[X_{1},X_{2},...,X_{n}]^{\intercal}$ is defined as
@@
where $h_{i}$ and $J_{ij}$ are the Lagrange multiplies to meet the first- and second-order cross-moment constraints, respectively. The detailed derivation is presented in A. Since our goal is to model the joint distribution that outputs the probability for any specified configuration, we drop the subscript $I$ in Eq. (7) and replace $P_{I}$ by $p(\boldsymbol{x})$:
@@
$=\frac{1}{Z}\exp{\left(\boldsymbol{h}\cdot\boldsymbol{x}+\frac{1}{2}\boldsymbol{x^{T}}\boldsymbol{J}\boldsymbol{x}\right)}\,,$
where $\boldsymbol{x}=\left[x_{1},x_{2},\dots,x_{n}\right]^{\intercal}$, $\boldsymbol{h}=\left[h_{1},h_{2},\dots,h_{n}\right]^{\intercal}$, and
@@
Notice that Eq. (9) and Eq. (7) are equivalent, because any configuration of $\boldsymbol{x}$ is indexed by $I$. Eq. (9) is the Ising model for describing ferromagnetism, where $h_{i}$ and $J_{ij}$ can be interpreted as the external magnetic field acting on each component/spin and the pairwise coupling effects between each spin pair, respectively [28, 29]. While the classic Ising model features a sparse $\boldsymbol{J}$ because it considers only the pairwise interactions between neighboring spins on a lattice, Eq. (9) in this study represents a long-range or fully-connected Ising model, accounting for interactions between all pairs of structures in a region.
@@ 3.2 Parameter estimation
The Boltzmann machine learning, a widely used method for the inverse Ising problem, obtains the maximum likelihood estimation of the Ising model parameters through gradient-based optimization. The log-likelihood $\mathcal{L}$ of the parameters $\boldsymbol{h}$ and $\boldsymbol{J}$ given observations $\mathcal{D}=\{\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)},…,\boldsymbol{x}^{(M)}\}$ is given as:
@@
$=\sum_{i=1}^{n}h_{i}\left(\frac{1}{M}\sum_{m=1}^{M}\sum_{i=1}^{n}x^{(m)}_{i}\right)+\sum_{i>j}J_{ij}\left(\frac{1}{M}\sum_{m=1}^{M}\sum_{i>j}x^{(m)}_{i}x^{(m)}_{j}\right)-\ln{Z(\boldsymbol{h},\boldsymbol{J})}$
@@
where $\mathcal{M}_{1,i}^{\mathcal{D}}$ and $\mathcal{M}_{2,ij}^{\mathcal{D}}$ denote the first- and second-order cross-moments calculated from the observations $\mathcal{D}$. The final line of Eq. (11) reveals that the likelihood function hinges on $\mathcal{M}_{1,i}^{\mathcal{D}}$ and $\mathcal{M}_{2,ij}^{\mathcal{D}}$ rather than on the random samples in $\mathcal{D}$. This suggests that the Ising model can be applied when $\mathcal{M}_{1,i}^{\mathcal{D}}$ and $\mathcal{M}_{2,ij}^{\mathcal{D}}$ are derived from sources other than random samples. Consequently, the Ising model can be constructed using fragility functions [36, 37, 38], analytic models [39, 40, 41], or a combination of both [42], not requiring information beyond the current PBEE practice.
@@
Finally, we obtain the updating rule for the maximum likelihood estimation:
@@
where $\alpha$ is the learning rate of the algorithm. Owing to the convexity of the inverse Ising problem, Eq .(13) is guaranteed to converge to a unique solution [18]. However, the exact evaluation of $\mathcal{M}_{1,i}$ and $\mathcal{M}_{2,ij}$ at each step requires calculating Eqs. (5) and (6) over $2^{n}$ configurations, which is generally infeasible. Hence, Monte Carlo sampling techniques are often used to approximate $\mathcal{M}_{1,i}$ and $\mathcal{M}_{2,ij}$ [43]. In this study, we adopted Gibbs sampling due to its applicability to high-dimensional distributions and ease of implementation.
@@ 3.3 Mean-field approximation
In this subsection, we introduce the mean-field approximation, one of the simplest approximation methods [44, 45, 18] for the inverse Ising problem, to facilitate the interpretation of the Ising model parameters. By averaging the interactions between components, the mean value of each component for a specified $\boldsymbol{h}$ and $\boldsymbol{J}$ is given by the solution of the following self-consistent equation:
@@
The accuracy and validity of the mean-field approximation depend on the magnitude and homogeneity of pairwise interactions, as the method averages these interactions and neglects individual effects [46]. Although various methods have been proposed to improve the applicability of the mean-field approximation, such as the Thouless-Anderson-Palmer (TAP) method [44, 18] and the Sessak-Monasson expansion [47], we found that none of these methods achieve satisfactory accuracy in regional seismic analysis. Therefore, our motivation of using the mean-field approximation is to facilitate a qualitative interpretation of the model parameters.

### v2

@@ Abstract
This study introduces the long-range Ising model from statistical mechanics to the Performance-Based Earthquake Engineering (PBEE) framework for regional seismic damage analysis. The application of the PBEE framework at a regional scale involves estimating the damage states of numerous structures, typically performed using fragility function-based stochastic simulations. However, these simulations often assume conditional independence or employ simplistic dependency models among the damage states of structures, leading to significant misrepresentation of regional risk. The Ising model addresses this issue by converting the available information on binary damage states (safe or failure) into a joint probability mass function, leveraging the principle of maximum entropy. The Ising model offers two main benefits: (1) it requires only the first- and second-order cross-moments, enabling seamless integration with the existing PBEE framework, and (2) it provides meaningful physical interpretations of the model parameters, facilitating the uncovering of insights not apparent from data. To demonstrate the proposed method, we applied the Ising model to $156$ buildings in Antakya, Turkey, using post-hazard damage evaluation data, and to $182$ buildings in Pacific Heights, San Francisco, using simulated data from the Regional Resilience Determination (R2D) tool. In both instances, the Ising model accurately reproduces the provided information and generates meaningful insights into regional damage. The study also investigates the change in Ising model parameters under varying earthquake magnitudes, along with the mean-field approximation, further facilitating the applicability of the proposed approach.
@@ 1 Introduction
The Performance-Based Earthquake Engineering (PBEE) framework has been extensively developed and utilized for over 20 years, providing a systematic method for risk-informed seismic design and post-hazard decision-making [1, 2]. To incorporate the various uncertainties that propagate from ground motions to the seismic performance of a structure, the PBEE framework considers four groups of random variables: (1) ground motion intensity measure (IM), (2) engineering demand parameter (EDP), (3) damage measure or damage state (DS), and (4) loss or decision variable (DV). Following the convention of probability theory, we use uppercase letters to represent random variables (e.g., DS) and lowercase letters to represent outcomes of random variables (e.g., ds). Given an earthquake scenario, IM is typically described as a probability density function $f(im)$, or as a mean annual probability of exceedance $\lambda(im)$, assuming a Poisson process model for the temporal occurrence of earthquakes. This intensity measure model is followed by the conditional distribution of EDP given IM, denoted by $f(edp|im)$. Subsequently, the conditional distribution of DS given EDP, $f(ds|edp)$, is established, which can also be expressed as a complementary cumulative distribution function $G(ds|edp)$. Finally, $f(ds|edp)$ is followed by the distribution of DV given DS, denoted by $f(dv|ds)$. To obtain the exceedance probability $\lambda(dv)$ of a decision variable, the probability density/mass functions $f(\cdot)$ are often transformed to complementary cumulative distribution functions $G(\cdot)$, yielding the integral:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\lambda(dv)=\int_{ds}\int_{edp}\int_{im}G(dv|ds)|d\,G(ds|edp)||d\,G(edp|im)||d\,\lambda(im)|\,.}}$ (1)
@@
Recently, the application of the PBEE framework has been actively extended to a regional scale [7, 8, 9, 10]. In this context, Eq. (1) is typically solved using Monte Carlo methods that simulate regional responses to hazards, often referred to as regional seismic simulations. This simulation approach is favored because a system-level decision variable may depend on the damage states of all components, rendering Eq. (1) high-dimensional. Similar to the component-level applications, a regional-scale PBEE analysis can be decomposed into four modules: (1) IM generation, (2) EDP estimation, (3) DS evaluation, and (4) DV assessment. For each of these modules, physics-based or statistical methods/models can be employed, offering a spectrum of variations. For instance, to generate IM samples, one can simulate seismic wave propagation [11] to create synthetic seismograms or employ empirical ground motion prediction equations [12]. Nonlinear structural dynamic analysis can be performed to generate EDP samples [10], while several regression models directly connecting EDP to IM are also available [13, 14]. The DS can be obtained from structural analysis or inferred from IM or EDP samples using DS-IM or DS-EDP fragility curves. Lastly, DV can be assessed using loss curves that relate DS or EDP to DV [15], or physics-based simulations [9].
Physics-based simulations can accurately capture the complex interactions and correlations among structural responses and functionalities. However, they are often restricted to scenario-based analyses due to their extensive computational costs. On the other hand, statistical approaches, such as fragility function-based simulations of DS given EDP or IM [16], are computationally feasible but require many ad-hoc, sometimes subjective, assumptions about statistical models. Specifically, applying structural-level fragility functions to regional-scale analysis entails constructing a fragility field that summarizes the joint information across the region. To capture the spatial dependency among structural damage, the fragility field demands additional assumptions and parameters than individual fragility curves. The common practice of using individual fragility functions under the assumption of conditional independence (conditional on site-specific intensity measures) can significantly underestimate system-level failure probabilities [17].
@@
This paper is organized as follows. Section 2 overviews the current approaches for evaluating structural damage in the regional-scale PBEE. Subsequently, Section 3 introduces the formulation and parameter estimation of the Ising model in the context of regional seismic analysis. This section also develops the mean-field approximation to qualitatively explain the Ising model parameters. Section 4 offers interpretations of the Ising model parameters in the language of regional seismic risk analysis and explores their relationship with the mean and correlation coefficients of the damage states of structures. Section 5 demonstrates the validity and engineering relevance of the Ising model through two examples using real-world post-hazard survey data and synthetic simulation data, emphasizing the insights derived from the Ising model. Finally, concluding remarks and future research directions are presented in Section 6.
@@ 2 Current approaches for probabilistic structural damage predictions on a regional-scale
For an individual structure, the following fragility function is commonly used to estimate the probability of exceeding a damage state:
@@
The missing information in the above procedure is the correlation between damage states contributed by the similarities of structures. Bazzurro and Luco Nicolas [22] emphasized that accounting for the similarity of structures represents a key distinction between analyzing a single structure and a portfolio of structures in the seismic loss estimation. Similarities between structures, such as the number of stories, construction materials, and structural type, are particularly pronounced in residential neighborhoods where structures share similar designs and are built following the same codes. This contributes to the correlation between their damage states [17], in addition to the contribution from the spatial coherency of ground motions. Other than the damage states, the importance of accurately modeling the correlations between EDPs of structures was emphasized by Kang et al. [23], who proposed a framework to account for the correlations between EDP residuals, defined as the residual term in a regression function relating EDP to IM. This approach was later applied to regional seismic loss assessment [24, 25]. Hamidia et al. [26] and Asjodi et al. [27] also investigated the correlation between EDPs stemming from the similarities of mechanical properties, employing computer vision-based methods.
To model the structure-to-structure damage correlation, Lee and Kiremidjian [28] presented a simplified approach by assuming equi-correlation among all structures. The study also concluded that the challenge of modeling the damage correlation lies in approximating the joint discrete damage distribution. Due to the difficulty of defining a joint distribution for discrete damage states, Baker and Cornell [29] proposed using the joint continuous distribution for the decision variables given the engineering demand parameters, $f(\boldsymbol{dv}|\boldsymbol{edp})$, by integrating $f(\boldsymbol{dm}|\boldsymbol{edp})$ and $f(\boldsymbol{dv}|\boldsymbol{dm})$, thereby bypassing the need for a joint distribution of discrete damage states. An alternative approach, as detailed in [30, 31], models fragility using continuous demand and capacity. This method becomes popular because of its convenience.
@@ 3 Long-range Ising model as a joint probability mass function for damage states
The Ising model was originally conceived as a mathematical model for ferromagnetism in statistical mechanics [33, 34]. A key feature of the model is that it represents the maximum entropy distribution constrained by the means and pairwise correlations, making it applicable to a wide variety of fields beyond statistical mechanics [35, 36, 18]. Provided with the first- and second-order cross-moments of the component states in a system, the Ising model is the least structured joint probability distribution for the component states that reproduces the given information. In this section, we first formulate the Ising model within the context of regional seismic analysis, followed by a discussion on its parameter estimation. Finally, we present the mean-field approximation, which offers a qualitative understanding of the relationship between the Ising model parameters and the given statistics.
@@ 3.1 Formulation
Consider a region consisting of $n$ structures under an earthquake and assume that each structure can be in one of the binary states of safe or failure. Here, safe and failure are relative terms with respect to a target performance level. The group of structures has $N=2^{n}$ configurations, where the probability of the $I$-th configuration is expressed as:
@@
The entropy $S$ for $\boldsymbol{X}=[X_{1},X_{2},...,X_{n}]^{\intercal}$ is defined as:
@@
where $h_{i}$ and $J_{ij}$ are the Lagrange multipliers to meet the first- and second-order cross-moment constraints, respectively. The detailed derivation is presented in A. Since our goal is to model the joint distribution that outputs the probability for any specified configuration, we drop the subscript $I$ in Eq. (7) and replace $P_{I}$ by $p(\boldsymbol{x})$:
@@
$=\frac{1}{Z}\exp{\left(\boldsymbol{h}^{\intercal}\boldsymbol{x}+\frac{1}{2}\boldsymbol{x}^{\intercal}\boldsymbol{J}\boldsymbol{x}\right)}\,,$
where $\boldsymbol{x}=\left[x_{1},x_{2},\dots,x_{n}\right]^{\intercal}$, $\boldsymbol{h}=\left[h_{1},h_{2},\dots,h_{n}\right]^{\intercal}$, and:
@@
Notice that Eq. (9) and Eq. (7) are equivalent, because any configuration of $\boldsymbol{x}$ is indexed by $I$. Eq. (9) is the Ising model for describing ferromagnetism, where $h_{i}$ and $J_{ij}$ can be interpreted as the external magnetic field acting on each component/spin and the pairwise coupling effects between each spin pair, respectively [33, 34]. The classic Ising model features a sparse $\boldsymbol{J}$ because it considers only the pairwise coupling effects between neighboring spins on a lattice. However, in this study, Eq. (9) represents a long-range or fully-connected Ising model, accounting for interactions between all pairs of structures in a region.
@@ 3.2 Parameter estimation
The Boltzmann machine learning, a widely used method for the inverse Ising problem, obtains the maximum likelihood estimation of the Ising model parameters through gradient-based optimization. The log-likelihood $\mathcal{L}$ of the parameters $\boldsymbol{h}$ and $\boldsymbol{J}$ given a collection of $M$ observations, $\mathcal{D}=\{\boldsymbol{x}^{(1)},\boldsymbol{x}^{(2)},\ldots,\boldsymbol{x}^{(M)}\}$, is expressed as:
@@
$=\sum_{i=1}^{n}h_{i}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\frac{1}{M}\sum_{m=1}^{M}x^{(m)}_{i}}}\right)+\sum_{i>j}J_{ij}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\frac{1}{M}\sum_{m=1}^{M}x^{(m)}_{i}x^{(m)}_{j}}}\right)-\ln{Z(\boldsymbol{h},\boldsymbol{J})}$
@@
where $\mathcal{M}_{1,i}^{\mathcal{D}}$ and $\mathcal{M}_{2,ij}^{\mathcal{D}}$ denote the first- and second-order cross-moments calculated from the observations $\mathcal{D}$. The final line of Eq. (11) reveals that the likelihood function hinges on $\mathcal{M}_{1,i}^{\mathcal{D}}$ and $\mathcal{M}_{2,ij}^{\mathcal{D}}$ rather than on the random samples in $\mathcal{D}$. This suggests that the Ising model can be applied when $\mathcal{M}_{1,i}^{\mathcal{D}}$ and $\mathcal{M}_{2,ij}^{\mathcal{D}}$ are derived from sources other than random samples. Consequently, the Ising model can also be constructed using fragility functions developed either from empirical data [41, 42, 43], analytical models [44, 45, 46], or a combination of both [47], without requiring information beyond the current PBEE practice.
@@
Finally, we obtain the gradient descent updating rule for the $k$-th step of the maximum likelihood estimation:
@@
where $\alpha$ is the learning rate of the algorithm. The initial values are set to $h_{i}^{(1)}=\tanh^{-1}\mathcal{M}_{1,i}$ and $J_{ij}^{(1)}=0$, based on the mean-field solution with zero pairwise interaction, which will be discussed in the following subsection. Owing to the convexity of the inverse Ising problem, Eq. (13) is guaranteed to converge to a unique solution, i.e., a unique set of Ising model parameters that reproduces the provided information [18]. However, the exact evaluation of $\mathcal{M}_{1,i}$ and $\mathcal{M}_{2,ij}$ at each step requires calculating Eqs. (5) and (6) over $2^{n}$ configurations, which is generally infeasible. Hence, Monte Carlo sampling techniques are often used to approximate $\mathcal{M}_{1,i}$ and $\mathcal{M}_{2,ij}$ [48]. In this study, we adopted Gibbs sampling due to its applicability to high-dimensional distributions and ease of implementation.
@@ 3.3 Mean-field approximation
In this subsection, we introduce the mean-field approximation, one of the simplest approximation methods [49, 50, 51, 18] for the inverse Ising problem, to facilitate the interpretation of the Ising model parameters. By averaging the interactions between components, the mean value of each component for a specified $\boldsymbol{h}$ and $\boldsymbol{J}$ is given by the solution of the following self-consistent equation:
@@
The accuracy and validity of the mean-field approximation depend on the magnitude and homogeneity of the pairwise interactions, as this method averages these interactions and neglects individual effects. Notably, the mean-field solution becomes exact when the damage states are independent, i.e., for all pairs, $J_{ij}=0$ Although various methods have been proposed to improve the applicability of the mean-field approximation, such as the Thouless-Anderson-Palmer (TAP) method [49, 18] and the Sessak-Monasson expansion [52], we found that none of these methods achieve satisfactory accuracy in regional seismic analysis with non-negligible correlations. Therefore, our usage of the mean-field approximation is to facilitate a qualitative interpretation of the model parameters.